OpenStudy (anonymous):

solve using the elimination method: 5r-8s=-25 8r+5s=49

OpenStudy (anonymous):

it seems if you try to cancel out one variable it ends up cancelling out the other one as well this means there are no solutions for this system of equations

OpenStudy (anonymous):

5r - 8s = -25 -----------(1) 8r + 5s = 49 -----------(2) Multiply (1) by 8 and (2) by 5 40r - 64s = - 200 --------(3) 40r + 25s = 245 --------(4) subtracting (4) from (1), we get -89s = - 445 s = -445/-89 s = 5 substitute s = 5 in (1) 5r - 8(5) = -25 5r - 40 = -25 5r = -25 + 40 5r = 15 r = 15/5 r = 3 so solution set = (3,5)

OpenStudy (anonymous):

so jm92 was wrong........ this pair of linear equations has a solution if u r satisfied with the answer, pls click GOOD ANSWER button for me....☺